A motorcycle traveling 50 mph overtakes a car traveling at 40 mph that had a 3 hr. Head start. How far from the starting point are the two vehicles
Answer by stanbon(75887)

A motorcycle traveling 50 mph overtakes a car traveling at 40 mph that had a 3 hr. Head start. How far from the starting point are the two vehicles
-------
Bike DATA:
rate = 50 mph ; time = x hrs ; distance = r*t = 50x miles
-----
Car DATA:
rate = 40 mph ; time = x+3 hrs ; distance = 40x+120 miles
------
Equation:
distance = distance
50x = 40x+120
---
10x = 120
x = 12 hrs
----
Ans: distance = 50*12 = 600 miles
-------------
Cheers,
Stan H.
-------------